OpenStudy (anonymous):

Coco's coffee charges different prices based on the number of pounds purchased. The pricing scale is modeled by the function below, where w is the weight in pounds purchased. Orders placed directly through the Web site are discounted by 1/3, but a shipping fee of $2.50 is added. Write a pricing function for orders placed through the Web site.

OpenStudy (anonymous):

Someone please help

OpenStudy (anonymous):

I'm thinking ._.

OpenStudy (anonymous):

So I have never done this math before but most math comes naturally to me so here is my answer: Let t represent the total amount of pounds purchased. Let a represent the total amount of money spent on pounds purchased. t / (1/3) + $2.50 = a

OpenStudy (anonymous):

I also just pulled this out of the air so I am most likely wrong. I also might be wright and if I am I consider myself a genius XD. Is my answer one of your possible answers?

OpenStudy (anonymous):

These are what the options look like. It doesn't make sense to me

OpenStudy (anonymous):

Woe you go to Lakeland or is it just a cowincadence the is a girl that goes to lakeland with you name and has to use the same program "brain honey" we use for comp class. BTW I also am completely lost with this math ;_; Im only a 9th grader >.>

OpenStudy (kash_thesmartguy):

\[\frac{ 1 }{ 3 }w+2.50=Total Charge\]

OpenStudy (anonymous):

I'm still so confused xD And no, i have no idea what Lakeland is. How would 1/3w +2.50 get me the answer?